Find the ordered pair that is a solution to the graphs y=2x + 3, y=x-2
Answer by stanbon(75887)

Find the ordered pair that is a solution to the graphs y=2x + 3, y=x-2
Equation:
2x+3 = x-2
x = -5
-----
Since y = x-2, y = -7
=========================
Cheers,
Stan H.
=========================